OpenStudy (anonymous):

Need help finding the derivative of: y = x / sqrt(1-x^2)

OpenStudy (anonymous):

quotient rule

OpenStudy (anonymous):

Did it but Im messing up somewhere

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

1 / (1-x^2)^(3/2)

OpenStudy (anonymous):

let's try it a different way... \[\frac{ d }{ dx }\left( \frac{ x }{ \sqrt{1-x^2} } \right)=\frac{ d }{ dx }x\left( \sqrt{1-x^2}\right)^{-1} \]\[= 1\cdot \left( \sqrt{1-x^2}\right)^{-1} + x \cdot (-1)\left( \sqrt{1-x^2}\right)^{-2}\cdot (-2x)\]

OpenStudy (anonymous):

Hold on, just working through it

OpenStudy (anonymous):

don't be scared of the quotient rule you probably need it over one denominator in any case, especially if you have to do something with the derivative once you find it

OpenStudy (anonymous):

alright so 2x^2 in the numerator then combine the two sqrt(1-x^2) quantities in the denom?

OpenStudy (anonymous):

\[=\frac{2x^2+\sqrt{1-x^2}}{1-x^2}\]

OpenStudy (anonymous):

sat - I did the quotient rule but I think i messed up when pulling out like terms

OpenStudy (anonymous):

don't forget that you need a common denominator

OpenStudy (anonymous):

pg can i draw out what i did and you check the work?

OpenStudy (anonymous):

shoot, i screwed up because it's a sqrt. should have written it using fractional exponents. sorry, give me a sec to redo

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

Use the quotient rule.. I understand it completely just dont know where i went wrong

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left( \frac{ x }{ \sqrt{1-x^2} } \right)=\frac{ d }{ dx }x\cdot (1-x^2)^{-\frac{ 1 }{ 2 }}\]\[=(1-x^2)^{-\frac{ 1 }{ 2 }}+x\cdot \left(-\frac{ 1 }{ 2 }\right)(1-x^2)^{-\frac{ 3 }{ 2 }}(-2x)\]

OpenStudy (anonymous):